PROBLEM SOLVING.

A. SET. Let A, B and C are sets and U be universal set.

U = {-1, 0, 1, 2, 3, 4, 5, 6, a, b, c, d, e}

A = {-1, 1, 2, 4}

B = {0, 2, 4, 6}

C = {b, c, d}

Find for the following. Show complete solutions.

1. 𝐡 βˆͺ 𝐢

2. 𝐴 βˆ’ 𝐡 π‘₯ 𝐢

3. π‘ƒπ‘œπ‘€π‘’π‘Ÿ 𝑠𝑒𝑑 π‘œπ‘“ 𝐢

4. |𝑃(𝐡)|

B. SEQUENCES. Consider the sequence {Sn} defined by Sn = 2n – 5, where 𝒏 β‰₯ βˆ’πŸ.

Find for:

1. βˆ‘1𝑖=βˆ’1 𝑆𝑖 (5 pts)

C. RELATION. Consider X = {-3, -2, -1, 0, 1} defined by (x,y) ∈ R if x β‰₯ y.

Find for:

1. Elements of R

2. Domain and Range of R

3. Draw the digraph

4. Identify the properties of R


1
Expert's answer
2021-04-27T01:44:59-0400

Solution:

(A):

U = {-1, 0, 1, 2, 3, 4, 5, 6, a, b, c, d, e}

A = {-1, 1, 2, 4}

B = {0, 2, 4, 6}

C = {b, c, d}

1. 𝐡 βˆͺ 𝐢 = {0, 2, 4, 6, b, c, d}

2. "𝐵 \times 𝐶=\{(0,b),(0,c),(0,d),(2,b),(2,c),(2,d),(4,b),(4,c),(4,d),(6,b),(6,c),(6,d)\}"

Now, "A-B\times C=\{-1, 1, 2, 4\}"

3. π‘ƒπ‘œπ‘€π‘’π‘Ÿ 𝑠𝑒𝑑 π‘œπ‘“ 𝐢"=\{\phi,\{b\}, \{c\},\{d\},\{b,c\},\{c,d\},\{b,d\},\{b,c,d\}\}"

4. |𝑃(𝐡)|"=2^n" , where n is the number of elements in set B.

"|P(B)|=2^4=16"

(B):

"S_n=2n-5,n\ge-1"

"\Sigma_{-1}^1 S_i=S_{-1}+S_0+S_1
\\=2(-1)-5+2(0)-5+2(1)-5
\\=-2+0+2-15
\\=-15"

(C):

Consider X = {-3, -2, -1, 0, 1} defined by (x,y) ∈ R if xΒ β‰₯Β y.

1.

"R=\{(-3,-3),(-2,-2),(-1,-1),(0,0),(1,1),(-2,-3),(-1,-3),\\(0,-3),(1,-3),(-1,-2),(0,-2),(1,-2),(0,-1),(1,-1)(1,0)\}"

2. Domain of R"=\{-3,-2,-1,0,1\}"

And range of R"=\{-3,-2,-1,0,1\}"

3. Digraph of R:



4.

Reflexive:

It is clearly reflexive as "(a,a)\in R, \forall a\in X"

Symmetric:

It is clearly not symmetric as "a\ge b" but "b\ge a" is not true, "\forall a,b \in X"

Moreover, "(0,-2)\in R" but "(-2,0)\not\in R"

Transitive:

"a\ge b, b\ge c \Rightarrow a\ge c" which is true "\forall a,b,c \in X"

Hence, it is transitive.